a locomotive of 60 tonne mass pull some carriages of total mass 240 tonne, a distance of half a kilometre up a 1 in 200 incline with an acceleration of 0.2ms^-2. if the overall rolling resistance is 5% of a normal reaction, what is the total work done?


1
Expert's answer
2022-02-17T04:36:37-0500

Given:

"m=\rm 60\: t+240\: t=300\: t=3*10^5\: kg"

"d=500\:\rm m"

"\sin\theta=1/200"

"a=0.2\:\rm m/s^2"

"\mu=0.05"


The net force

"F=ma=3*10^5*0.2=6*10^4\:\rm N"

The total work done

"W=Fd=6*10^4\:\rm N*500\: m=3*10^7\:\rm J"